# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ E.5.1. Hardness of $\text{IVM}^{\pm}[Q_{\text{nh}}]$
E.5.2. Hardness of $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$
We reduce the OuMv problem to $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$. The encoding of the matrix and the vectors is exactly as in Section E.5.1. The only difference is how we obtain the result of $\bm{u}_{r}^{\top}\bm{M}\bm{v}_{r}$ in each round $r\in[n]$. We first execute the necessary updates to encode $\bm{u}_{r}$ in $R$. Then, we delete the contents of $T$. Afterwards, we insert, one by one, each value $j$ with $\bm{v}_{r}(j)=1$ to $T$. After each insert, we request the enumeration of the delta in the result of $Q_{\text{nh}}$ and check whether the delta contains at lease one tuple. The main observation is that $\bm{u}^{\top}\bm{M}\bm{v}=1$ if and only if one of the enumeration requests in round $r$ reports at least one tuple. The overall number of updates is as in Section E.5.1. The additional cost caused by each enumeration request is constant.
E.5.3. Hardness for Arbitrary Non-Hierarchical Queries

### v3

@@ F.5.1. Hardness of $\text{IVM}^{\pm}[Q_{\text{nh}}]$
F.5.2. Hardness of $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$
We reduce the OuMv problem to $\text{IVM}^{\pm}_{\delta}[Q_{\text{nh}}]$. The encoding of the matrix and the vectors is exactly as in Section F.5.1. The only difference is how we obtain the result of $\bm{u}_{r}^{\top}\bm{M}\bm{v}_{r}$ in each round $r\in[n]$. We first execute the necessary updates to encode $\bm{u}_{r}$ in $R$. Then, we delete the contents of $T$. Afterwards, we insert, one by one, each value $j$ with $\bm{v}_{r}(j)=1$ to $T$. After each insert, we request the enumeration of the delta in the result of $Q_{\text{nh}}$ and check whether the delta contains at lease one tuple. The main observation is that $\bm{u}^{\top}\bm{M}\bm{v}=1$ if and only if one of the enumeration requests in round $r$ reports at least one tuple. The overall number of updates is as in Section F.5.1. The additional cost caused by each enumeration request is constant.
F.5.3. Hardness for Arbitrary Non-Hierarchical Queries